The first month of the winter in Bloomington, is a cold month, with an average

temperature ranging between max. 5.3 °C and min. -4.1 °C. During this time

there are a lot of snows and the roads are covered with ice. Which chemical

compound(s) you use to lower the freezing point of ice for the streets of

Bloomington and why? Suppose, you dissolve 87.0 g of rock salt in 500.0 g of

water. What will be the freezing point of that solution? (K f = 1.86 o C/molal for

water).


1
Expert's answer
2021-01-21T03:48:29-0500

"to - tf = \frac{w}{m}\times\frac{1000}{Wg}\times Kf"

"0 - tf = \frac{87}{58.5}\times\frac{1000}{500}\times 1.86"

"tf = -5.53"0C

By applying salt to an icy road, the melting point of the ice is decreased, and the ice will melt more quickly, making driving safer. Sodium chloride (NaCl) and either calcium chloride (CaCl 2 ) or magnesium chloride (MgCl 2 ) are used most frequently, either alone or in a mixture. Sodium chloride is the least expensive option, but is less effective because it only dissociates into two ions instead of three.